# Balanced cross-Kerr coupling for superconducting qubit readout

### v1

@@ Abstract
Dispersive readout, the standard method for measuring superconducting qubits, is limited by multiphoton qubit-resonator processes arising even at moderate drive powers. These processes degrade performance, causing dispersive readout to lag behind single- and two-qubit gates in both speed and fidelity. In this work, we propose a novel readout method, termed “junction readout”. Junction readout leverages the nonperturbative cross-Kerr interaction resulting from coupling a qubit and a resonator via a Josephson junction. Furthermore, by adding a capacitive coupling in parallel to the junction, Purcell decay can be suppressed without the need for a Purcell filter. We also show that junction readout is more robust against deleterious multiphoton processes, and offers greater flexibility for resonator frequency allocation. Crucially, junction readout achieves superior performance compared to dispersive readout while maintaining similar hardware overhead. Numerical simulations show that junction readout can achieve fidelities exceeding $99.99\%$ in under $30$ ns, making it a promising alternative for superconducting qubit readout with current hardware.
@@
Fast and high-fidelity qubit measurement is a cornerstone of quantum information processing and fault-tolerant quantum computing. For example, in quantum error correction (QEC) protocols, each round of QEC relies on rapid single-shot readout of ancilla qubits to detect errors. Recent breakthroughs achieving break-even performance in surface [1, 2] and bosonic codes [3, 4, 5] have relied on dispersive readout, a standard tool for measuring superconducting qubits. This readout operates by introducing a qubit-state dependent shift in the resonator frequency, enabling qubit state inference without directly disturbing the qubit [6, 7, 8].
@@
Despite the improvements in readout fidelity and integration times [9, 10, 11, 12, 13], dispersive readout still lags behind the performance of the best single- and two-qubit gates. Increasing the readout drive power to enhance fidelity and speed often leads to measurement-induced transitions [14, 15, 16, 17, 18, 19, 20, 21] which are detrimental to error correction protocols as they introduce correlated errors [22, 23]. Resetting such states, which typically involve $\sim$ 5 to 10 photons [19], is challenging even with the use of leakage reduction techniques [24, 25, 26, 27]. Thus, achieving fast, high-fidelity readout with low leakage rates remains an open problem.
To further complicate matters, state-of-the-art dispersive readout often requires a Purcell filter to prevent the qubit from decaying through the readout channel [28, 29]. While effective, adding a Purcell filter increases the readout system’s footprint, complicates calibration, and makes multiplexing more challenging. These challenges have spurred interest in intrinsically Purcell-protected qubits and readout methods as compact, scalable alternatives for next-generation quantum processors [30, 31, 32, 33, 34].
Here, we propose an approach to mediate a nonperturbative dispersive qubit-resonator interaction, enabling high-fidelity and fast measurement without the need for a Purcell filter. Even when accounting for reduced readout efficiency and finite qubit lifetime, our approach achieves an order-of-magnitude improvement over state-of-the-art readout systems with comparable hardware overhead.
Theory of junction readout.—
In circuit QED, the dispersive interaction $\sum_{i_{t}}\chi_{i_{t}}\ket{i_{t}}\bra{i_{t}}\hat{a}^{\dagger}\hat{a}$, with |$\ket{i_{t}}$ a bare transmon state and $\hat{a}$ the resonator’s annihilation operator, is usually realized by capacitively coupling a qubit and a resonator that are widely detuned in frequency [8]. Here, we seek an alternative circuit that mediates this interaction, with the interaction strength $2\chi_{z}=\chi_{1_{t}}-\chi_{0_{t}}$ engineered to be nonperturbative (i.e., independent of the qubit-resonator detuning and not obtained after some Schrieffer-Wolff-type frame transformation).
@@
Alternative strategies for engineering a nonperturbative Kerr interaction have been explored [32, 31, 33, 35, 36, 34]. In contrast to these approaches, which often involve complex circuit designs, rely on special symmetries, or encode qubits in spatially delocalized modes, our proposed circuit, shown in Fig. 1, offers a simpler solution with hardware overhead comparable to that of dispersive readout. In this circuit, the transmon (green) is coupled to a readout resonator (blue) through a Josephson junction in parallel with a capacitor whose role will be discussed below. The Hamiltonian of this circuit takes the form
@@
$\begin{split}\hat{H}&=4E_{c}(\hat{n}_{t}-n_{g})^{2}-E_{J}\cos\hat{\varphi}_{t}+\omega_{r}\hat{a}^{\dagger}\hat{a}\\ &\quad-E_{J_{c}}\cos(\hat{\varphi}_{t}-\hat{\varphi}_{r})+J\hat{n}_{t}\hat{n}_{r}\\ &\equiv\hat{H}_{\text{tr}}+\hat{H}_{\text{r}}+\hat{H}_{\text{int}},\end{split}$ (1)
@@
where $\hat{\varphi}_{t}$ and $\hat{n}_{t}$ are the transmon’s phase and charge operators, and $E_{J}$ and $E_{c}$ are its Josephson and charging energies, respectively. The resonator has frequency $\omega_{r}$, and its phase operator is given by $\hat{\varphi}_{r}=\varphi_{\mathrm{zpfr}}(\hat{a}+\hat{a}^{\dagger})$ where $\varphi_{{\rm zpfr}}=(2\pi/\Phi_{0})\sqrt{\hbar Z_{r}/2}$ is the phase zero-point fluctuations. Furthermore, $E_{Jc}$ is the Josephson energy of the coupling junction, and $J$ is the capacitive coupling strength between the transmon and the resonator. The gate charge $n_{g}$ is included explicitly, as it has been shown to affect the onset of measurement-induced transitions [37, 18, 19, 21]. As we show below, our proposed circuit ensures a high measurement critical photon number regardless of the value of the gate charge.
@@
To second order in phase fluctuations, the cos-cos interaction of the first term leads to $\hat{\varphi}_{t}^{2}\hat{\varphi}_{r}^{2}$, thereby mediating a nonperturbative cross-Kerr coupling. As discussed in Ref. [38], this results in large measurement critical photon numbers and no Purcell decay. When supplemented with the appropriate resonator drive [38], this interaction emulates the longitudinal coupling discussed in Ref. [32], enabling fast and accurate discrimination of the pointer states. Here, we show that the circuit of Fig. 1 achieves the aforementioned benefits with a significantly simplified design. On the other hand, to first order in phase fluctuations, the sin-sin interaction term is an unwanted Jaynes-Cummings interaction that causes multiphoton resonances [15, 16, 17, 18, 19], thereby lowering the critical photon number, see Sec. S3.1 of [39]. A potential mitigation strategy involves increasing the detuning between the qubit and resonator while increasing the qubit-resonator coupling to maintain a constant dispersive shift [40, 41]. However, this requires the use of low-frequency qubits strongly coupled to high-frequency resonators which poses its own set of challenges. Additionally, for moderate detunings, this approach leads to Purcell-limited qubit lifetimes, necessitating the use of a Purcell filter.
Instead, here we propose to cancel this unwanted sin-sin term by introducing a parallel capacitance to the Josephson junction which mediates a charge-charge coupling $J\hat{n}_{t}\hat{n}_{r}$, see Fig. 1. The coupling strength $J$ is specifically chosen to cancel the Purcell decay by setting the $0\leftrightarrow 1$ matrix element of the interaction Hamiltonian to zero, as expressed by the condition
$\begin{split}-&E_{J_{c}}\bra{1_{t},0_{r}}\sin\hat{\varphi}_{t}\sin\hat{\varphi}_{r}\ket{0_{t},1_{r}}\\ &+J\bra{1_{t},0_{r}}\hat{n}_{r}\hat{n}_{t}\ket{0_{t},1_{r}}=0,\end{split}$ (3)
where |$\ket{j_{r}}$ denotes a bare resonator state. Intuitively, this condition describes destructive interference between the current paths through the capacitor and the junction, flowing from the qubit to the readout feedline. Importantly, since the cancellation condition involves only the computational subspace, the optimal coupling strength $J$ is insensitive to fluctuations in the gate charge $n_{g}$. Moreover, as shown in Fig. 1(c), the residual Purcell decay remains minimal even when the condition Eq. 3 is not perfectly satisfied. For instance, with a $\sim 10\%$ imperfection in the junction or capacitor fabrication, the Purcell decay time $T_{1}$ remains on the order of $\sim 1\ \text{ms}$ for the chosen parameters [39], highlighting the robustness of this scheme to fabrication errors.
Figure 2: (a) Cross-Kerr coupling $\chi_{z}$ and (b) resonator self-Kerr $K_{r}$ between the transmon and the resonator for varying resonator impedance $Z_{r}$ and coupling junction energy $E_{Jc}$. The contour lines indicate lines of constant (a) cross-Kerr ranging from $-2$ to $-10\>\textrm{MHz}$ and (b) self-Kerr ranging from $-100$ to $-800\>\textrm{kHz}$. The star marks the parameter used in the readout simulations of Fig. 4.
@@
We now turn to a discussion of the optimal choice of parameters. To achieve fast readout we aim for a large transmon-resonator cross-Kerr interaction. However, as is evident from Eq. 2, a nonperturbative cross-Kerr term (resulting from $\hat{\varphi}_{t}^{2}\hat{\varphi}_{r}^{2}$) is inherently accompanied by a nonperturbative self-Kerr nonlinearity on the resonator (resulting from $\hat{\varphi}_{r}^{4}$). A large self-Kerr can bananize the coherent state in the resonator [42, 43], potentially reducing readout fidelity by hindering the clear separation of pointer states and limiting the maximum photon population in the resonator. Moreover, this distortion renders the conventional linear measurement filter suboptimal for state assignment [44, 45, 46].
@@
To achieve a large cross-Kerr interaction without compromising the linearity of the readout resonator, we optimize the resonator impedance $Z_{r}$. From Eq. 2, the leading contribution of the cross-Kerr strength is given by $\chi_{z}\simeq-\varphi_{\mathrm{zpfr}}^{2}E_{Jc}\sqrt{2E_{C}/E_{J,\rm{total}}}/2$, with $E_{J,\rm{total}}=E_{J}+E_{Jc}$ which we keep constant. The self-Kerr nonlinearity of the resonator inherited from the coupling junction is approximately $K_{r}\simeq-E_{Jc}\varphi_{\mathrm{zpfr}}^{4}/4.$ As a result, the cross-Kerr interaction decreases quadratically with the resonator phase zero-point fluctuations $\varphi_{{\rm zpfr}}$, while the self-Kerr decreases quartically with $\varphi_{{\rm zpfr}}$. Conversely, both $\chi_{z}$ and $K_{r}$ increase linearly with respect to the coupling junction energy $E_{Jc}$. Since $\varphi_{{\rm zpfr}}$ depends on the resonator impedance $Z_{r}$, we optimize $Z_{r}$ and $E_{Jc}$ to achieve a large cross-Kerr coupling with minimal self-Kerr nonlinearity. Fig. 2(a) and (b) show the cross-Kerr coupling strength $\chi_{z}$ and the resonator’s self-Kerr $K_{r}$ extracted from exact numerical diagonalization of Eq. 1 for a range of resonator impedances $Z_{r}$ and coupling strengths $E_{Jc}$. For realistic circuit parameters, we are able to achieve a cross-Kerr coupling comparable to standard dispersive readout, typically ranging from 2 to 10 MHz. Furthermore, by slightly reducing the resonator impedance below the conventional $50~{}\Omega$— something that is easily achievable experimentally— the self-Kerr nonlinearity can be tuned to match, or in some cases even fall below, the typical values of 100 to 500 kHz observed in standard dispersive readout.
@@
Interestingly, we also observe that, in junction readout, the dispersive shift shows minimal variation with increasing resonator photon numbers, in contrast to dispersive readout, where the dispersive shift decreases significantly. This difference arises from the nature of the qubit-resonator interaction which is different in the two schemes. Indeed, in junction readout, the cos-cos interaction leads to smaller higher-order corrections to the dispersive shift at large photon numbers compared to dispersive readout, enabling a faster readout as the cross-Kerr coupling remains large even when the resonator is populated, see [39] for further details.
@@
The above analysis reveals that junction readout produces qubit-state-dependent frequency shifts similar to those of dispersive readout, suggesting comparable performance at first glance. In the following, we highlight the significant advantages offered by junction readout, demonstrating its superiority over dispersive readout in key aspects, namely its larger robustness against measurement-induced transitions, as well as faster and higher fidelity measurements.
@@
The quantum nondemolition (QND) nature of dispersive readout is challenged by multiphoton processes arising due to accidental degeneracies between the qubit and the resonator [15]. This phenomenon, also referred to as ionization, occurs when two states of the transmon-resonator system become resonant as photons populate the resonator, leading to a sudden population transfer from the resonator to the transmon  [15, 19, 16, 18, 17, 21, 47]. Because this process can typically involve highly-excited states of the transmon which are charge sensitive, the resonator critical photon number $n_{\rm{crit}}$ at which these multiphoton processes occur can fluctuate widely with gate charge [19, 48, 18, 37]. We now show junction readout is far more robust against such multiphoton processes than dispersive readout.
To assess this robustness, we compute the critical photon numbers using branch analysis, a numerical tool predicting the onset of ionization and that has been shown to match experimental observations [48]; see Ref. [39] for further details in the context of junction readout. Fig. 3 shows, given an arbitrary gate charge in the range [0,0.5], the percentage of instances where the critical photon number exceeds $n_{\rm{crit}}=35$ for both junction readout (blue) and dispersive readout (red), across a wide resonator frequency range of $\omega_{r}/2\pi\sim 6-10\>\textrm{GHz}$. To ensure a fair comparison between both approaches, the dispersive shift is fixed to $\chi_{z}/2\pi\simeq 9\>\textrm{MHz}$ across all values of $\omega_{r}$. For the dispersive readout, this requires adjusting the qubit-resonator coupling at all values of the resonator frequency [8]. The value of $n_{\rm{crit}}$ reported in Fig. 3 is arbitrary but chosen to correspond to a readout fidelity of approximately 99.99%, given the dispersive shift used here, see Sec. S3.2 of [39]. We first observe that for junction readout there exists a wide $\sim 3\>\textrm{GHz}$ range of resonator frequencies for which junction readout reaches near $100\>\%$ certainty of achieving a critical photon number exceeding $35$ photons. In contrast, for dispersive readout there exists no frequency range for which the critical photon number consistently exceeds this threshold. Furthermore, dispersive readout has less flexibility for frequency allocation as the dispersive shift depends on the detuning, and often requires the readout resonator to be close in frequency to the qubit for a large dispersive shift. On the other hand, junction readout relies on a nonperturbative cross-Kerr interaction which does not depend on the detuning, and thus offers greater flexibility for frequency allocation and potentially can alleviate frequency crowding issues for large-scale chips, whilst maintaining a large cross-Kerr for fast readout.
Figure 3: Percentage of instances where the critical photon number exceeds $n_{\rm{crit}}=35$ with varying gate charge on the transmon for both junction readout (blue) and dispersive readout (red). In both cases, the dispersive shift is set to a constant value of $\chi_{z}/2\pi\simeq 9\>\textrm{MHz}$. With a qubit frequency of $\omega_{q}/2\pi\simeq 5.68\>\textrm{GHz}$, the shaded region corresponds to resonator frequencies that are outside of the dispersive regime.
Another important consequence of the cancellation condition of Eq. 3 is to push ionization to larger photon numbers. Indeed, from perturbation theory, when the transmon ionizes to an excited state outside of the computational manifold, the transition occurs through virtual excitations, sequentially climbing the intermediate states between the initial state and the final ionized state [48]. The cancellation condition in Eq. 3, however, eliminates the matrix element responsible for the transition between the ground state and the first excited state, thereby preventing the initial step of leaving the computational subspace from the ground state. Additionally, while not exact, this cancellation approximately suppresses the matrix element responsible for transitions from the first to the second excited state, further reducing ionization when the qubit begins in its first excited state.
@@
Figure 4: (a) IQ-plane of 11,200 single-shot heterodyne readout simulations of a transmon with junction readout. The black dashed line is the optimal discriminator of the two blobs corresponding to the qubit prepared in the ground (blue) or excited state (red). (b) Histogram of the integrated signal using an optimal discriminator. For both (a) and (b) the integrated time is $t_{m}\simeq 20\>\textrm{ns}$. (c) Assignment error obtained from the stochastic heterodyne readout simulations compared to coherent state approximated assignment error. Using the coherent state approximated assignment error, we also show the assignment error for readout efficiency of $\eta=0.5$ (where $0\leq\eta\leq 1$) as well as when $T_{1}$ is $30$ or $120\>\mu s$. Moreover, we show that quantum optimal control (QOC) further improves the readout fidelity. Here, the transmon charging energy is $E_{C}/2\pi=300\>\textrm{MHz}$ with $E_{J,\rm{total}}/E_{C}=50$ and gate charge $n_{g}=0.0$. The resonator frequency is $\omega_{r}/2\pi=9.375\>\textrm{GHz}$, with resonator impedance $Z_{r}=40\>\Omega$. The junction coupling strength is $E_{Jc}/2\pi=8\>\textrm{GHz}$. This set of parameter results in a cross-Kerr of $\chi_{z}/2\pi\simeq 10\>\textrm{MHz}$, and a critical photon number of $n_{\rm{crit}}=65$.

### v2

@@ Abstract
Dispersive readout, the standard method for measuring superconducting qubits, is limited by multiphoton qubit-resonator processes arising even at moderate drive powers. These processes degrade performance, causing dispersive readout to lag behind single- and two-qubit gates in both speed and fidelity. In this work, we propose a novel readout method, termed “junction readout”. Junction readout leverages the nonperturbative cross-Kerr interaction resulting from coupling a qubit and a resonator via a Josephson junction. Furthermore, by adding a capacitive coupling in parallel to the junction, Purcell decay induced by the exchange coupling can be suppressed. We also show that junction readout is more robust against deleterious multiphoton processes, and offers greater flexibility for resonator frequency allocation. Crucially, junction readout achieves superior performance compared to dispersive readout while maintaining similar hardware overhead. Numerical simulations show that junction readout can achieve fidelity exceeding $99.99\%$ in under $30$ ns of integration time, making it a promising alternative for superconducting qubit readout with current hardware.
@@
Fast and high-fidelity qubit measurement is a cornerstone of quantum information processing and fault-tolerant quantum computing. For example, in quantum error correction (QEC) protocols, each round of QEC relies on rapid single-shot readout of ancilla qubits to detect errors. Recent breakthroughs achieving break-even performance in surface [GoogleQuantumAI:2023, GoogleQuantumAI:2024] and bosonic codes [Ofek:2016, Sivak:2023, Brock:2025] have relied on dispersive readout, a standard tool for measuring superconducting qubits. This readout operates by introducing a qubit-state dependent shift in the resonator frequency, enabling qubit state inference without directly disturbing the qubit [Blais_PRA, Wallraff_disp_readout, RMP].
Despite the improvements in readout fidelity and integration times [McClure2016, Walter_disp_readout, Sunada_intrinsic, swiadek2023enhancing, Spring:2024], dispersive readout still lags behind the performance of the best single- and two-qubit gates. Increasing the readout drive power to enhance fidelity and speed often leads to measurement-induced state transitions [Johnson2012, MIST_1, Lescanne:2019, Dynamics_of_transmon_ionzation, Reminiscence_chaos, MIST_2, hazra_2024, Dumas2024] which are detrimental to error correction protocols as they introduce correlated errors [Varbanov:2020, Miao:2023]. Resetting such states, which typically involve $\sim$ 5 to 10 photons [MIST_2], is challenging even with the use of leakage reduction techniques [McEwen:2021, Battistel:2021, Marques:2023, Lacroix:2023]. Thus, achieving fast, high-fidelity readout with low leakage rates remains an open problem.
Figure 1: Junction readout circuit (a) with and (b) without flux loop or low-frequency mode. A transmon (green) is coupled to the readout resonator (blue) through a Josephson junction (red) and a capacitor. Unlike in (a), in (b) the transmon nonlinearity is entirely inherited from the coupling junction, see Ref. [Supplementary_Information]. (c) Jaynes–Cummings induced Purcell lifetime $T_{1}^{\mathrm{Purcell}}$ of the transmon for varying capacitive coupling strengths $J$. The black dashed line indicates where the cancellation condition of Eq. 3 is met. There, $J/2\pi\simeq 32.8\>\textrm{MHz}$, corresponding to a coupling capacitance around $10\>\textrm{fF}$. The qubit and resonator frequencies are $\omega_{q}/2\pi=5.672$ GHz and $\omega_{r}/2\pi=9.375$ GHz, respectively, and the resonator decay rate is $\kappa/2\pi=8$ MHz, see Sec S2 of [Supplementary_Information].
To further complicate matters, state-of-the-art dispersive readout often requires a Purcell filter to prevent the qubit from decaying through the readout channel [Reed2010, Jeffrey_purcell]. While effective, adding a Purcell filter increases the readout system’s footprint, complicates calibration, and makes multiplexing more challenging. These challenges have spurred interest in intrinsically Purcell-protected qubits and readout methods as compact, scalable alternatives for next-generation quantum processors [dimon_Gambetta, Diniz_CrossKerr, Didier_longitudinal, Remy_cross_kerr, Pmon].
Here, we propose an approach to mediate a nonperturbative dispersive qubit-resonator interaction, enabling high-fidelity and fast measurement while suppressing the dominant channel contributing to Purcell decay. Even when accounting for reduced readout efficiency and finite qubit lifetime, our approach achieves an order-of-magnitude improvement over state-of-the-art readout systems with comparable hardware overhead.
Theory of junction readout.—
@@
Figure 2: (a) Cross-Kerr coupling $\chi_{z}$ and (b) resonator self-Kerr $K_{r}$ between the transmon and the resonator for varying resonator impedance $Z_{r}$ and coupling junction energy $E_{Jc}$. The contour lines indicate lines of constant (a) cross-Kerr ranging from 2-2 to $10\>\textrm{MHz}$ and (b) self-Kerr ranging from 100-100 to $800\>\textrm{kHz}$. The star marks the parameter used in the readout simulations of Fig. 4.
@@
Alternative strategies for engineering a nonperturbative Kerr interaction have been explored [Didier_longitudinal, Diniz_CrossKerr, Remy_cross_kerr, Quarton_ye_2021, ye2024ultrafast, Pmon]. In contrast to these approaches, which often involve complex circuit designs, rely on special symmetries, or encode qubits in spatially delocalized modes, our proposed circuits, shown in Fig. 1, offer a simpler solution with hardware overhead comparable to that of dispersive readout. In the circuit Fig. 1 (a), the transmon (green) is coupled to a readout resonator (blue) through a Josephson junction in parallel with a capacitor whose role will be discussed below. The Hamiltonian of this circuit takes the form
@@
$\begin{split}\hat{H}&=4E_{C}(\hat{n}_{t}-n_{g})^{2}-E_{J}\cos\hat{\varphi}_{t}\\ &\quad+\omega_{r}\hat{a}^{\dagger}\hat{a}\\ &\quad-E_{J_{c}}\cos(\hat{\varphi}_{t}-\hat{\varphi}_{r})+J\hat{n}_{t}\hat{n}_{r}\\ &\equiv\hat{H}_{\text{tr}}+\hat{H}_{\text{r}}+\hat{H}_{\text{int}},\end{split}$ (1)
@@
where the first (second) line corresponds to the transmon (resonator) Hamiltonian $\hat{H}_{\text{tr}}$ ($\hat{H}_{\text{r}}$) while the last line corresponds to the interaction Hamiltonian $\hat{H}_{\text{int}}$. Furthermore, $\hat{\varphi}_{t}$ and $\hat{n}_{t}$ are the transmon’s phase and charge operators which satisfy the relation $e^{i\hat{\varphi}_{t}},\hat{n}_{t}]=-e^{i\hat{\varphi}_{t}}$, and $E_{J}$ and $E_{C}$ are its Josephson and charging energies, respectively. The resonator has frequency $\omega_{r}$, and its phase operator is given by $\hat{\varphi}_{r}=\varphi_{\mathrm{zpfr}}(\hat{a}+\hat{a}^{\dagger})$ where $\varphi_{{\rm zpfr}}=(2\pi/\Phi_{0})\sqrt{\hbar Z_{r}/2}$ is the phase zero-point fluctuations. Furthermore, $E_{Jc}$ is the Josephson energy of the coupling junction, and $J$ is the capacitive coupling strength between the transmon and the resonator. The gate charge $n_{g}$ is included explicitly, as it has been shown to affect the onset of measurement-induced state transitions [Malekakhlagh:2022, Reminiscence_chaos, MIST_2, Dumas2024]. As we show below, our proposed circuit ensures a high measurement critical photon number regardless of the value of the gate charge.
@@
To second order in phase fluctuations, the cos-cos interaction of the first term leads to $\hat{\varphi}_{t}^{2}\hat{\varphi}_{r}^{2}$, thereby mediating a nonperturbative cross-Kerr coupling. As discussed in Ref. [chapple:2024], this results in large measurement critical photon numbers and no Purcell decay. When supplemented with the appropriate resonator drive [chapple:2024], this interaction emulates the longitudinal coupling discussed in Ref. [Didier_longitudinal], enabling fast and accurate discrimination of the pointer states. Here, we show that the circuit of Fig. 1 achieves the aforementioned benefits with a significantly simplified design. On the other hand, to first order in phase fluctuations, the sin-sin interaction term is an unwanted Jaynes-Cummings interaction that causes multiphoton resonances [MIST_1, Lescanne:2019, Dynamics_of_transmon_ionzation, Reminiscence_chaos, MIST_2], thereby lowering the critical photon number, see Sec S3A of [Supplementary_Information]. A potential mitigation strategy is to substantially increase the qubit–resonator detuning—well beyond what it is used in standard dispersive readout—while simultaneously increasing the coupling to keep the dispersive shift fixed [kurilovich:2025, connolly2025, dai2025]. In practice, this places the qubit at a small fraction of the resonator frequency (e.g., $\omega_{r}/\omega_{q}\sim 10$) and thus relies on low-frequency qubits strongly coupled to high-frequency resonators, which introduces its own set of challenges.
Instead, here we propose to cancel this unwanted sin-sin term by introducing a parallel capacitance to the Josephson junction which mediates a charge-charge coupling $J\hat{n}_{t}\hat{n}_{r}$, see Fig. 1. The coupling strength $J$ is specifically chosen to cancel the exchange (Jaynes–Cummings) coupling by setting the $0\leftrightarrow 1$ matrix element of the interaction Hamiltonian to zero, as expressed by the condition
[…]
@@
We now turn to a discussion of the optimal choice of parameters. To achieve fast readout we aim for a large transmon-resonator cross-Kerr interaction. However, as is evident from Eq. 2, a nonperturbative cross-Kerr term (resulting from $\hat{\varphi}_{t}^{2}\hat{\varphi}_{r}^{2}$) is inherently accompanied by a nonperturbative self-Kerr nonlinearity on the resonator (resulting from $\hat{\varphi}_{r}^{4}$). Resonator self-Kerr can distort the coherent state in the resonator [Sivak2019, Boutin2017], potentially reducing readout fidelity by hindering the clear separation of pointer states and limiting the maximum photon population in the resonator. Moreover, this distortion renders the conventional linear measurement filter suboptimal for state assignment [Bultink2018, Gembetta_optimal_readout, Danjou:2014].
@@
To achieve a large cross-Kerr interaction ($|\chi_{z}|/2\pi\sim 2-10$ MHz) without compromising the linearity of the readout resonator, we optimize the resonator impedance $Z_{r}$. From Eq. 2, the leading contribution of the cross-Kerr strength is given by $\chi_{z}\simeq-\varphi_{\mathrm{zpfr}}^{2}E_{Jc}\sqrt{2E_{C}/E_{J,\rm{total}}}/2$, with $E_{J,\rm{total}}=E_{J}+E_{Jc}$ which we keep constant. The self-Kerr nonlinearity of the resonator inherited from the coupling junction is approximately $K_{r}\simeq-E_{Jc}\varphi_{\mathrm{zpfr}}^{4}/2.$ As a result, the cross-Kerr interaction decreases quadratically with the resonator phase zero-point fluctuations $\varphi_{{\rm zpfr}}$, while the self-Kerr decreases quartically with $\varphi_{{\rm zpfr}}$. Conversely, both $\chi_{z}$ and $K_{r}$ increase linearly with respect to the coupling junction energy $E_{Jc}$. Since $\varphi_{{\rm zpfr}}$ depends on the resonator impedance $Z_{r}$, we optimize $Z_{r}$ and $E_{Jc}$ to achieve a large cross-Kerr coupling with minimal self-Kerr nonlinearity ($|K_{r}|/2\pi\lesssim 500$ kHz). Fig. 2(a) and (b) show the cross-Kerr coupling strength $\chi_{z}$ and the resonator’s self-Kerr $K_{r}$ extracted from exact numerical diagonalization of Eq. 1 for a range of resonator impedances $Z_{r}$ and coupling strengths $E_{Jc}$. For realistic circuit parameters, we are able to achieve a cross-Kerr coupling comparable to standard dispersive readout, typically ranging from 2 to 10 MHz. Furthermore, by slightly reducing the resonator impedance below $50~\Omega$—something that is easily achievable experimentally—the self-Kerr nonlinearity can be tuned to match, or in some cases even fall below, the typical values of 100 to 500 kHz observed in standard dispersive readout.
@@
Interestingly, we also observe that, in junction readout, the dispersive shift shows minimal variation with increasing resonator photon numbers, in contrast to dispersive readout, where the dispersive shift decreases significantly. This difference arises from the nature of the qubit-resonator interaction which is different in the two schemes. Indeed, in junction readout, the cos-cos interaction leads to smaller higher-order corrections to the dispersive shift at large photon numbers compared to dispersive readout, enabling a faster readout as the cross-Kerr coupling remains large even when the resonator is populated, see [Supplementary_Information] for further details.
@@
The above analysis reveals that junction readout produces qubit-state-dependent frequency shifts similar to those of dispersive readout, suggesting comparable performance at first glance. In the following, we highlight the significant advantages offered by junction readout, demonstrating its superiority over dispersive readout in key aspects, namely its larger robustness against measurement-induced state transitions, as well as faster and higher fidelity measurements.
@@
The quantum nondemolition (QND) nature of dispersive readout is challenged by multiphoton processes arising due to accidental degeneracies between the qubit and the resonator [MIST_1]. This phenomenon, also referred to as ionization, occurs when two states of the transmon-resonator system become resonant as photons populate the resonator, leading to a sudden population transfer from the resonator to the transmon  [MIST_1, MIST_2, Lescanne:2019, Reminiscence_chaos, Dynamics_of_transmon_ionzation, Dumas2024, Xiao2023]. Because this process can typically involve highly excited states of the transmon which are charge sensitive, the resonator critical photon number $n_{\rm{crit}}$ at which these multiphoton processes occur can fluctuate widely with gate charge [MIST_2, Dumas2024, Reminiscence_chaos, Malekakhlagh:2022]. Note that $n_{\rm{crit}}$ here differs from the Jaynes-Cummings $n_{\rm{crit}}^{jc}=(\Delta/2g)^{2}$ which marks the breakdown of the dispersive approximation in dispersive readout [RMP]. We now show junction readout is far more robust against such multiphoton processes than dispersive readout.
Figure 3: (a) Critical photon number averaged over gate charge $n_{g}\in[0,0.5]$ as a function of resonator frequency for junction readout (blue) and dispersive readout (red). In both cases, the dispersive shift is fixed at $|\chi_{z}|/2\pi\simeq 9$ MHz. Dashed horizontal line indicates $n_{\rm{crit}}=30$. (b) Minimum critical photon number over $n_{g}\in[0,0.5]$. Inset: ratio of the minimum critical photon numbers between junction and dispersive readout. Dashed horizontal line indicates where the ratio is $1$.
@@
Figure 4: (a) IQ-plane of 11,200 single-shot heterodyne readout simulations of a transmon with junction readout. The black dashed line is the optimal discriminator of the two blobs corresponding to the qubit prepared in the ground (blue) or excited state (red). (b) Histogram of the integrated signal using an optimal discriminator. For both (a) and (b) the integrated time is $t_{m}\simeq 20\>\textrm{ns}$. (c) Assignment error obtained from the stochastic heterodyne readout simulations compared to coherent state approximated assignment error. Using the coherent state approximated assignment error, we also show the assignment error for readout efficiency of $\eta=0.5$ (where $0\leq\eta\leq 1$) as well as when $T_{1}$ is $30$ or $120\>\mu s$. Moreover, we show that quantum optimal control (QOC) further improves the readout fidelity. Here, the transmon charging energy is $E_{C}/2\pi=300\>\textrm{MHz}$ with $E_{J,\rm{total}}/E_{C}=50$ and gate charge $n_{g}=0.0$. The resonator frequency is $\omega_{r}/2\pi=9.375\>\textrm{GHz}$, with resonator impedance $Z_{r}=40\>\Omega$ and decay rate $\kappa/2\pi=2|\chi_{z}|\simeq 20$ MHz. The junction coupling strength is $E_{Jc}/2\pi=8\>\textrm{GHz}$. This set of parameter results in a cross-Kerr of $|\chi_{z}|/2\pi\simeq 10\>\textrm{MHz}$, and a critical photon number of $n_{\rm{crit}}=65$.
To assess this robustness, we compute the critical photon numbers using branch analysis, a numerical tool predicting the onset of ionization and that has been shown to match experimental observations [Dumas2024]; see Ref. [Supplementary_Information] for further details in the context of junction readout. Fig. 3(a) shows the critical photon number averaged over different realizations of the gate charge $n_{g}\in[0,0.5]$, while panel (b) shows the minimum critical photon number across those realizations, for both junction readout (blue) and dispersive readout (red), over a wide range of resonator frequencies $\omega_{r}/2\pi\sim 6-10\>\textrm{GHz}$. To ensure a fair comparison between both approaches, the dispersive shift is fixed to $|\chi_{z}|/2\pi\simeq 9\>\textrm{MHz}$ across all values of $\omega_{r}$. For the dispersive readout, this requires adjusting the qubit-resonator coupling at all values of the resonator frequency [RMP]. We observe that junction readout consistently yields a higher average critical photon number than dispersive readout over nearly the entire range of resonator frequencies considered. Moreover, we show in Ref. [Supplementary_Information] that for junction readout, there exists a broad region of resonator frequencies in which over $95\%$ of gate charge realizations yield a critical photon number exceeding 30 (dashed horizontal line in Fig. 3(a)). In contrast, no such frequency range exists for dispersive readout where a comparable fraction of realizations consistently surpass this threshold—see [Supplementary_Information] for further details. Crucially, even in the worst-case scenario, the minimum critical photon number over realizations of the gate charge is larger for junction readout than for dispersive readout in most cases, as shown in Fig. 3(b).
The markedly higher ionization critical photon numbers observed in junction readout can be largely attributed to the cancellation condition of Eq. 3. Indeed, from perturbation theory, when the transmon ionizes to an excited state outside of the computational manifold, the transition occurs through virtual excitations, sequentially climbing the intermediate states between the initial state and the final ionized state [Dumas2024]. The cancellation condition in Eq. 3, however, eliminates the matrix element responsible for the transition between the ground state and the first excited state, thereby preventing the initial step of leaving the computational subspace from the ground state. Additionally, while not exact, this cancellation approximately suppresses the matrix element responsible for transitions from the first to the second excited state, further reducing ionization when the qubit begins in its first excited state, see Sec S3A of [Supplementary_Information] for more details.
Finally, we further note that dispersive readout has less flexibility for frequency allocation as the dispersive shift depends on the resonator frequency, and often requires the readout resonator to be close in frequency to the qubit for a large dispersive shift. On the other hand, junction readout relies on a nonperturbative cross-Kerr interaction which does not depend on the resonator frequency, and thus offers greater flexibility for frequency allocation and potentially can alleviate frequency crowding issues for large-scale chips, whilst maintaining a large cross-Kerr for fast readout.